A candy factory classifies its candies as (Carmel-5 types) (chocolate 7 types) ( Dark chocolate 10 types). A customer wants (4 types of Carmel) (5 types of Chocolate) (7 types of Dark Chocolate) and i need to know how many assortments are possible?
Answer by khwang(438)

Choose 4 among 5 Carmel types, there are C(5,4) = 5 =(5!/4!) possibilities
Choose 5 among 7 chocolate types, there are C(7,5) = (7!/(5!2!))
= 7*6/2 = 21 possibilities
Choose 7 among 10 Dark chocolate types, there are C(10,7) = (10!/(7!3!))
= 10*9*8/6 = 120 possibilities
5 + 21 + 120 = 146
Totally,there are 146 different assortments